# Understanding the Drag Torque in Common Envelope Evolution

### v1

@@ Abstract
Common envelope (CE) evolution is largely governed by the drag torque applied on the inspiralling stellar components by the envelope. Previous work has shown that idealized models of the torque based on a single body moving in rectilinear motion through an unperturbed atmosphere can be highly inaccurate. Progress requires new models for the torque that account for binarity. Toward this end we perform a new 3D global hydrodynamic CE simulation with the mass of the companion point particle set equal to the mass of the asymptotic giant branch star core particle to maximize symmetry and facilitate interpretation. First, we find that a region around the particles of a scale comparable to their separation contributes essentially all of the torque. Second, the density pattern of the torque-dominating gas and, to an extent, this gas itself, is roughly in corotation with the binary. Third, approximating the spatial distribution of the torquing gas as a uniform-density prolate spheroid whose major axis resides in the orbital plane and lags the line joining the binary components by a constant phase angle reproduces the torque evolution remarkably well, analogous to studies of binary supermassive black holes. Fourth, we compare the torque measured in the simulation with the predictions of a model that assumes two weak point-mass perturbers undergoing circular motion in a uniform background without gas self-gravity, and find remarkable agreement with our results if the background density is taken to be equal to a fixed fraction ($\approx 0.44$) of the density at the spheroid surface. Overall, this work makes progress toward developing simple time-dependent models of the CE phase.
@@ 1 INTRODUCTION
When a giant star engulfs a much smaller companion, the shared envelope drags on the companion and core, causing them to inspiral. This is known as a common envelope (CE) event and is important for a wide range of phenomena (see Ivanova et al. 2020, Röpke & De Marco 2023, and Schneider et al. 2025 for recent reviews). For stellar companions, the drag torque is dominated by gas dynamical friction, which has traditionally been modeled using Bondi-Hoyle-Lyttleton (BHL) theory (Hoyle & Lyttleton, 1939; Bondi & Hoyle, 1944) and extensions thereof (Dodd & McCrea, 1952; Ostriker, 1999; MacLeod & Ramirez-Ruiz, 2015; MacLeod et al., 2017; De et al., 2020).
@@
However, such models are unable to explain the drastic reduction in the gravitational drag torque seen to occur after the first periastron passage, during the slow spiral-in phase (e.g. Chamandy et al., 2019a, 2020). [^1] We note that ‘slow spiral-in’ can have slightly different meanings in other works on CE evolution. This reduction happens partly because the binary separation becomes comparable to the BHL accretion radius, causing the core and companion to experience a thrust from one another’s high-density wakes that partially cancels the drag from their own wakes. This effect has been discussed by Kim et al. (2008) (hereafter K08) in the idealized case of circular motion of equal point masses whose affect on the background density $\rho_{0}$ is weak enough to be accurately described as a linear perturbation in the variable $\lambda=(\rho-\rho_{0})/\rho_{0}$. K08 built on earlier work that studied the drag on a single such “perturber” in a circular orbit (e.g. Kim & Kim, 2007). Such studies make various assumptions and idealizations, not least that the linear treatment is adequate, and thus it is not immediately clear to what extent they can be applied to the CE case. [^2] We note that Sánchez-Salcedo & Brandenburg (2001) studied the drag force in numerical simulations that solve a more general version of the equations and allow the orbit of the point mass to evolve. However, these models (or perhaps extensions thereof) could potentially be extremely useful, so exploring the gap between such idealized models and global CE simulations may be a promising avenue. This is indeed one of the main goals of the present work (see Section 5.2).
@@
The scenario of a binary embedded in a gaseous medium also arises in other contexts, most famously that of binary supermassive black holes (BSMBHs), where gas dynamical friction may help to solve the “last parsec problem” (Escala et al. 2004; hereafter E04, Escala et al. 2005; Dotti et al. 2007; Chapon et al. 2013; Tang et al. 2017; Li & Lai 2022; Williamson et al. 2022), and compact object orbital tightening prior to coalescence, where it may help to solve the “last astronomical unit problem” (Tagawa et al., 2018). In CE studies there seems to be an analogous “last solar radius problem” that prevents simulations from reaching final separations small enough to explain observations of post-CE binary systems (Iaconi & De Marco, 2019). This problem is likely caused, in part, by the torque evolution described above; long-duration, high-resolution CE simulations find that the orbit, does, in fact, continue to decay due to dynamical friction on long timescales (Gagnier & Pejcha, 2023; Chamandy et al., 2024). While it has been shown that extra physics like ionization and MHD can play an important role during the CE phase (e.g. Ivanova et al., 2013; Sand et al., 2020), such effects are unlikely to have an order unity effect on the inspiral (e.g. Reichardt et al., 2020; Ondratschek et al., 2022; Chamandy et al., 2024). However, insufficient numerical resolution, including insufficiently small softening lengths, can cause the torque to be inaccurate, affecting the inspiral (Iaconi et al., 2018; Chamandy et al., 2024; Gagnier & Pejcha, 2024).
@@
Given the similarity between CE evolution and the other scenarios mentioned above, we explore how modelling approaches used in those cases might apply to the CE case. For analysing their simulation of inspiralling equal-mass BSMBHs, E04 developed a model for the dynamical friction torque that reproduces the evolution of the BSMBH separation at late times remarkably well, when the binary mass is much larger than the gas mass interior to the orbit. This condition is generally also met in the slow spiral-in phase of CE evolution. Their model takes advantage of the symmetry that arises when the components are of equal mass. They model the torque as that applied by a constant-density, constant-aspect ratio prolate spheroid concentric with the binary but with major axis (in the orbital plane) lagging the axis passing through the binary components by a constant phase angle. Inspired by E04, we try a similar approach for modelling the late-stage torque during CEE and for this purpose we run a CE simulation with companion mass equal to the core mass of the giant star.
@@
This paper is organized as follows. In Section 2 we describe the simulation methods. In Section 3, we summarize the idealized models to which our results are later compared, namely the lagging uniform density prolate spheroid model of E04 and the circularly orbiting double-perturber model of K08. In Section 4, we present results from our global CE simulation and compare the torque measured from the simulation with predictions of the above models. We discuss the successes and limitations of our work in Section 5. Finally, we summarize and conclude in Section 6.
@@ 2 METHODS
Our global hydrodynamic simulation is carried out with the 3D code astrobear (Cunningham et al., 2009; Carroll-Nellenback et al., 2013). The setup is identical to the ‘AGB’ run studied in Chamandy et al. (2020), except that the mass of the binary companion particle is made to equal that of the AGB core particle. The reader is referred to that paper for further details about the methods. The AGB primary has initial mass $M_{1}=1.78\,\mathrm{M_{\odot}}$ and radius $R_{1}=122.2\,\mathrm{R_{\odot}}$, and the binary is initialized in a circular orbit with separation $a_{\mathrm{i}}=124.0\,\mathrm{R_{\odot}}$. The initial AGB density and pressure profiles are taken from a 1D single-star simulation of a ZAMS $2\,\mathrm{M_{\odot}}$ star run using MESA (Paxton et al., 2011, 2013, 2015, 2019). Both the companion and core are modeled as gravitation-only point particles with mass $M_{2}=M_{\mathrm{1,c}}=0.534\,\mathrm{M_{\odot}}$. The particles are not sink particles, i.e., they have fixed mass and do not accrete. A spline function with a constant softening radius ($r_{\mathrm{soft}}$) of $2.41\,\mathrm{R_{\odot}}$ is used to smooth the particle potential (Hernquist & Katz, 1989; Springel, 2010). Five levels of adaptive mesh refinement are employed and the regions around the particles are resolved with a smallest cell size of $\delta_{5}=0.070\,\mathrm{R_{\odot}}$ out to about $12\,\mathrm{R_{\odot}}$ from each particle, resulting in the ratio $r_{\mathrm{soft}}/\delta_{5}=34.3$ cells per spline softening length. The central region of the initial primary is modified to account for the presence of the point particle (Ohlmann et al., 2017; Chamandy et al., 2018). The simulation domain is a cube of side $1150\,\mathrm{R_{\odot}}$, and the AGB star is placed at its centre, non-rotating in the lab frame. The initial density ($\rho_{\mathrm{amb}}$) and pressure ($P_{\mathrm{amb}}$) of the ambient medium are set to $1.0\times 10^{-9}\,{\rm g\,cm^{-3}}$ and $1.1\times 10^{4}\,{\rm dyn\,cm^{-2}}$, respectively. The simulation is stopped at $t=299\,{\rm d}$, when the total energy (which should ideally be conserved) has increased by $6\%$ (see appendix E of Chamandy et al. 2024 for a discussion about energy conservation in our CE simulations).
@@
[…]
where $\rho(\mathbf{s})$ is density of gas at position $\mathbf{s}$, $V$ is the volume of whole simulation box and $\mathbf{s}_{1}$ and $\mathbf{s}_{2}$ are the position vectors of the core and the companion, respectively. The $z-$component of the torque about the particle centre of mass is calculated using
$\tau_{z}=\frac{a}{m}\left(M_{2}{{\mbox{\boldmath$F$}}}_{1,\phi}+M_{\mathrm{1,c}}{{\mbox{\boldmath$F$}}}_{2,\phi}\right),$ (3)
where $a$ is the particle separation projected onto the $x$-$y$ ($r$-$\phi$) plane, which remains almost exactly parallel to the orbital plane of the particles during the simulation, and $m=M_{\mathrm{1,c}}+M_{2}$ is the combined mass of the binary point particles. In this work, $M_{\mathrm{1,c}}=M_{2}=m/2$. The separation $a$ and orbital evolution are shown in Figure 1. The quantity $-\tau_{z}$ (where the minus sign is introduced to cancel the minus sign resulting from net drag) is shown as a solid black line in both panels of Figure 2.
@@ 3.1 Uniform density lagging prolate spheroid model
The potential at a point $(x,y,z)$ inside an ellipsoid of uniform density $\overline{\rho}$ is given by
@@
where the coefficients $\alpha$, $\beta$, $\zeta$ and $\chi$ are given in Lamb (1879, p. 146-154, 700), Chandrasekhar (1969, p. 38-45), and Binney & Tremaine (2008, p. 83-95). Take the origin to be the particles’ orbital CM in the simulation and the $x$-axis to be the axis passing through both particles. For the special case of a constant-density prolate spheroid with major axis lagging this binary axis by the angle $\Delta\phi$, the $z$-component of torque at any point inside the spheroid is given by
@@
$\widetilde{A}$ and $\widetilde{B}$ are the semi-major and semi-minor axes of the spheroid, respectively, and $e$ is the spheroid eccentricity. The torque does not depend on the size of the ellipsoid because there is no force on the binary due to the ellipsoid mass located outside of the similar, concentric ellipsoid of the same orientation that passes through the particles. This is a consequence of Newton’s third theorem (Binney & Tremaine, 2008, p. 87). However, below we obtain $\overline{\rho}$ by averaging the density inside an ellipsoidal contour, so the overall size of the contour does play a role. In our choice of coordinate system, the expression for the aggregate torque on both particles reduces to (A)
@@ 3.2 Simplified double-perturber model
Another approach is to solve the governing equations after making various approximations. K08 solved numerically a linearized version of the hydrodynamics equations for two perturbers in a circular orbit, building on a study of a single perturber in a circular orbit Kim & Kim (2007), which in turn extended the work of Ostriker (1999) for rectilinear motion. The K08 model is valid if the perturbers are weak and if various complicating effects – gas self-gravity, orbital eccentricity, the density gradient and rotation of the gas in which the perturbers are embedded, the motion of the binary CM, and the time dependence due to orbital evolution – can all be safely neglected. Their model further assumes an ideal gas with adiabatic index $\gamma=5/3$, which is also assumed in our CE simulation.
Below we restrict our discussion to the case of equal mass perturbers. While K08 also focuses on this case, we note that their general model includes the mass ratio as a parameter. The force on each perturber can be written as
@@
where $\mathcal{I}_{R}{{\mbox{\boldmath$\hat{R}$}}}$ and $\mathcal{I}_{\phi}{{\mbox{\boldmath$\hat{\phi}$}}}$ characterize the radial and azimuthal components in the orbital plane,
@@
K08 (see also Kim & Kim 2007) solved numerically for the drag force exerted on the perturbers by their wakes and obtained fitting formulas that closely match the numerical solutions. The adiabatic sound speed is given by
@@
where $c_{\mathrm{s,0}}$ is the sound speed in the unperturbed medium. We shall see below that the subsonic regime, with $\mathcal{M}_{\mathrm{p}}<1$, is most relevant for this work. In this regime the K08 fitting formulas are
@@
An analytical solution has been derived by Desjacques et al. (2022) that agrees with the numerical solutions of K08, but as the analytical solution is cumbersome we choose to compare our solutions with the fitting formulas (17) and (18). [^3] See also Stahler (2010) for an alternative analytic solution to the problem.
The $z$-component of the torque on a single perturber in the rest frame of the CM of the perturbers is given by $-(a/2)\mathcal{F}\mathcal{I}_{\phi}$, where $\mathcal{F}$ is given by equation (12) and $\mathcal{I}_{\phi}$ by equation (14). Thus, for equal mass perturbers with equal Mach numbers, the torque on the binary is given by

### v3

@@ Abstract
Common envelope (CE) evolution is largely governed by the drag torque applied on the in-spiralling stellar components by the envelope. Previous work has shown that idealized models of the torque based on a single body moving in rectilinear motion through an unperturbed atmosphere can be highly inaccurate. Progress requires new models for the torque that account for binarity. Toward this end we perform a new 3D global hydrodynamic CE simulation with the mass of the companion point particle set equal to the mass of the asymptotic giant branch star core particle to maximize symmetry and facilitate interpretation. First, we find that a region around the particles of a scale comparable to their separation contributes essentially all of the torque. Second, the density pattern of the torque-dominating gas and, to an extent, this gas itself, is roughly in corotation with the binary. Third, approximating the spatial distribution of the torquing gas as a uniform-density prolate spheroid whose major axis resides in the orbital plane and lags the line joining the binary components by a constant phase angle reproduces the torque evolution remarkably well, analogous to studies of binary supermassive black holes. Fourth, we compare the torque measured in the simulation with the predictions of a model that assumes two weak point-mass perturbers undergoing circular motion in a uniform background without gas self-gravity, and find remarkable agreement with our results if the background density is taken to be equal to a fixed fraction ($\approx 0.44$) of the density at the spheroid surface. Overall, this work makes progress toward developing simple time-dependent models of the CE phase, for example by informing the development of drag force prescriptions for 1D spherically symmetric CE simulations, which could be used to explore the parameter space of luminous red novae or in binary population synthesis studies.
@@ 1 INTRODUCTION
When a giant star engulfs a much smaller companion, the shared envelope drags on the companion and core, causing them to inspiral. This is known as a common envelope (CE) event and is important for a wide range of phenomena (see Ivanova+20, Roepke+Demarco23, and Schneider+25 for recent reviews). For stellar companions, the drag torque is dominated by gas dynamical friction, which has traditionally been modeled using Bondi-Hoyle-Lyttleton (BHL) theory (Hoyle+Lyttleton39; Bondi+Hoyle44) and extensions thereof (Dodd+Mccrea52; Ostriker99; Macleod+Ramirez-ruiz15a; Macleod+17; De+20).
@@
However, such models are unable to explain the drastic reduction in the gravitational drag torque seen to occur after the first periastron passage, during the slow spiral-in phase (e.g. Chamandy+19b; Chamandy+20). [^1] We note that ‘slow spiral-in’ can have slightly different meanings in other works on CE evolution. This reduction happens partly because the binary separation becomes comparable to the BHL accretion radius, causing the core and companion to experience a thrust from one another’s high-density wakes that partially cancels the drag from their own wakes. This effect has been discussed by Kim+08 (hereafter Kim+08) in the idealized case of circular motion of equal point masses whose effect on the background density $\rho_{0}$ is weak enough to be accurately described as a linear perturbation in the variable $\lambda=(\rho-\rho_{0})/\rho_{0}$. Kim+08 built on earlier work that studied the drag on a single such “perturber” in a circular orbit (e.g. Kim+Kim07). Such studies make various assumptions and idealizations, not least that the linear treatment is adequate, and thus it is not immediately clear to what extent they can be applied to the CE case. [^2] We note that Sanchez-salcedo+Brandenburg01 studied the drag force in numerical simulations that solve a more general version of the equations and allow the orbit of the point mass to evolve. However, these models (or perhaps extensions thereof) could potentially be extremely useful, so exploring the gap between such idealized models and global CE simulations may be a promising avenue. This is indeed one of the main goals of the present work (see Section 5.2).
@@
The scenario of a binary embedded in a gaseous medium also arises in other contexts, most famously that of binary supermassive black holes (BSMBHs), where gas dynamical friction may help to solve the “last parsec problem” (Escala+04; hereafter Escala+04, Escala+05; Dotti+07; Chapon+13; Tang+17; Li+Lai22; Williamson+22), and compact object orbital tightening prior to coalescence, where it may help to solve the “last astronomical unit problem” (Tagawa+18). In CE studies there seems to be an analogous “last solar radius problem” that prevents simulations from reaching final separations small enough to explain observations of post-CE binary systems (Iaconi+Demarco19). This problem is likely caused, in part, by the torque evolution described above; long-duration, high-resolution CE simulations find that the orbit, does, in fact, continue to decay due to dynamical friction on long timescales (Gagnier+Pejcha23; Chamandy+24). While it has been shown that extra physics like ionization and MHD can play an important role during the CE phase (e.g. Ivanova+13a; Sand+20), such effects are unlikely to have an order unity effect on the inspiral (e.g. Reichardt+20; Ondratschek+22; Chamandy+24). However, insufficient numerical resolution, including insufficiently small softening lengths, can cause the torque to be inaccurate, affecting the inspiral (Iaconi+18; Chamandy+24; Gagnier+Pejcha25).
@@
Given the similarity between CE evolution and the other scenarios mentioned above, we explore how modelling approaches used in those cases might apply to the CE case. For analysing their simulation of inspiralling equal-mass BSMBHs, Escala+04 developed a model for the dynamical friction torque that reproduces the evolution of the BSMBH separation at late times remarkably well, when the binary mass is much larger than the gas mass interior to the orbit. This condition is generally also met in the slow spiral-in phase of CE evolution. Their model takes advantage of the symmetry that arises when the components are of equal mass. They model the torque as that applied by a constant-density, constant-aspect ratio prolate spheroid concentric with the binary but with major axis (in the orbital plane) lagging the axis passing through the binary components by a constant phase angle. Inspired by Escala+04, we try a similar approach for modelling the late-stage torque during CE evolution and for this purpose we run a CE simulation with the companion mass equal to the core mass of the giant star.
@@
This paper is organized as follows. In Section 2 we describe the simulation methods. In Section 3, we summarize the idealized models to which our results are later compared, namely the lagging uniform density prolate spheroid model of Escala+04 and the circularly orbiting double-perturber model of Kim+08. In Section 4, we present results from our global CE simulation and compare the torque measured from the simulation with predictions of the above models. We discuss the successes and limitations of our work in Section 5. Finally, we summarize and conclude in Section 6.
@@ 2 METHODS
Our global hydrodynamic simulation is carried out with the 3D code astrobear (Cunningham+09; Carroll-nellenback+13). The setup is identical to the ‘AGB’ run studied in Chamandy+20, except that the mass of the binary companion particle is made to equal that of the AGB core particle. The reader is referred to that paper for further details about the methods. The AGB primary has initial mass $M_{1}=1.78\,\mathrm{M_{\odot}}$ and radius $R_{1}=122.2\,\mathrm{R_{\odot}}$, and the binary is initialized in a circular orbit with separation $a_{\mathrm{i}}=124.0\,\mathrm{R_{\odot}}$. The initial AGB density and pressure profiles are taken from a 1D single-star simulation of a ZAMS $2\,\mathrm{M_{\odot}}$ star run using mesa (Paxton+11; Paxton+13; Paxton+15; Paxton+19). Both the companion and core are modeled as gravitation-only point particles with mass $M_{2}=M_{\mathrm{1,c}}=0.534\,\mathrm{M_{\odot}}$. The particles are not sink particles, i.e., they have fixed mass and do not accrete. A spline function with a constant softening radius ($r_{\mathrm{soft}}$) of $2.41\,\mathrm{R_{\odot}}$ is used to smooth the particle potential (Hernquist+Katz89; Springel10). The 1D stellar profile inside $r=r_{\mathrm{soft}}$ is modified to account for the presence of the AGB core particle by solving a modified Lane-Emden equation and iterating over the particle mass until the correct interior mass is recovered (Ohlmann+17; Chamandy+18). Five levels of adaptive mesh refinement are employed and the regions around the particles are resolved with a smallest cell size of $\delta_{5}=0.070\,\mathrm{R_{\odot}}$ out to about $12\,\mathrm{R_{\odot}}$ from each particle, resulting in the ratio $r_{\mathrm{soft}}/\delta_{5}=34.3$ cells per spline softening length. The simulation domain is a cube of side $1150\,\mathrm{R_{\odot}}$, and the AGB star is placed at its centre, non-rotating in the lab frame. The initial density ($\rho_{\mathrm{amb}}$) and pressure ($P_{\mathrm{amb}}$) of the ambient medium are set to $1.0\times 10^{-9}\,{\rm g\,cm^{-3}}$ and $1.1\times 10^{4}\,{\rm dyn\,cm^{-2}}$, respectively. The simulation is stopped at $t=299\,{\rm d}$, when the total energy (which should ideally be conserved) has increased by $6\%$ (see appendix E of Chamandy+24 for a discussion about energy conservation in our CE simulations).
@@
${{\mbox{$F$}}}_{1}=GM_{1,c}\int_{V}\rho({{\mbox{$s$}}})\frac{{{\mbox{$s$}}}-{{\mbox{$s$}}}_{1}}{|{{\mbox{$s$}}}-{{\mbox{$s$}}}_{1}|^{3}}\,\mathrm{d}^{3}{{\mbox{$s$}}},$ (1)
@@
where $\rho(\mathbf{s})$ is density of gas at position $\mathbf{s}$, $V$ is the volume of the whole simulation box and $\mathbf{s}_{1}$ and $\mathbf{s}_{2}$ are the position vectors of the core and the companion, respectively. The $z$-component of the torque about the particle centre of mass is calculated using
$\tau_{z}=\frac{a}{m}\left(M_{2}{{\mbox{$F$}}}_{1,\phi}+M_{\mathrm{1,c}}{{\mbox{$F$}}}_{2,\phi}\right),$ (3)
where $a$ is the particle separation projected onto the $x$-$y$ ($r$-$\phi$) plane, which remains almost exactly parallel to the orbital plane of the particles during the simulation, and $m=M_{\mathrm{1,c}}+M_{2}$ is the combined mass of the binary point particles. In this work, $M_{\mathrm{1,c}}=M_{2}=m/2$. The separation $a$ and orbital evolution are shown in Figure 1. The quantity $\tau_{z}$ (where the minus sign is introduced to cancel the minus sign resulting from net drag so that we can plot a positive quantity) is shown as a solid black line in both panels of Figure 2.
@@ 3 IDEALIZED MODELS FOR COMPARISON WITH SIMULATION RESULTS
Our first method of approximation is to model the gravitational torque on the particles as that exerted by a uniform ellipsoid rotating at the instantaneous orbital angular frequency, with its major axis situated in the orbital plane and lagging the line joining the particles by a phase angle.
@@ 3.1 Uniform density lagging prolate spheroid model
The gravitational potential at a point $(x,y,z)$ inside an ellipsoid of uniform density $\overline{\rho}$ is given by
@@
where the coefficients $\alpha$, $\beta$, $\zeta$ and $\chi$ are given in lamb1879, Chandrasekhar69, and Binney+Tremaine08. Take the origin to be the particles’ orbital centre of mass (CM) in the simulation and the $x$-axis to be the axis passing through both particles. For the special case of a constant-density prolate spheroid with major axis lagging this binary axis by the angle $\Delta\phi$, the $z$-component of torque at any point inside the spheroid is given by
@@
$\widetilde{A}$ and $\widetilde{B}$ are the semi-major and semi-minor axes of the spheroid, respectively, and $e$ is the spheroid eccentricity. As a consequence of Newton’s third theorem (Binney+Tremaine08, p. 87), the ellipsoid’s mass outside the similar, concentric ellipsoid (of the same orientation) that passes through the particles exerts no net force on the binary. Therefore, the torque does not depend directly on the size of the ellipsoid. However, below we obtain $\overline{\rho}$ by averaging the density inside an ellipsoidal contour, so the overall size of the contour does play a role. In our choice of coordinate system, the expression for the aggregate torque on both particles reduces to (A)
@@ 3.2 Simplified double-perturber model
Another approach is to solve the governing equations after making various approximations. Kim+08 solved numerically a linearized version of the hydrodynamics equations for two perturbers in a circular orbit, building on a study of a single perturber in a circular orbit, Kim+Kim07, which in turn extended the work of Ostriker99 for rectilinear motion. The Kim+08 model is valid if the perturbers are weak and if various complicating effects – gas self-gravity, orbital eccentricity, the density gradient and rotation of the gas in which the perturbers are embedded, the motion of the binary CM, and the time dependence due to orbital evolution – can all be safely neglected. Their model further assumes an ideal gas with adiabatic index $\gamma=5/3$, which is also assumed in our CE simulation.
Below we restrict our discussion to the case of equal mass perturbers. While Kim+08 also focuses on this case, we note that their general model includes the mass ratio as a parameter. The force on each perturber can be written as
@@
where $\mathcal{I}_{R}{{\mbox{$\hat{R}$}}}$ and $\mathcal{I}_{\phi}{{\mbox{$\hat{\phi}$}}}$ characterize the radial and azimuthal components in the orbital plane,
@@
Kim+08 (see also Kim+Kim07) solved numerically for the drag force exerted on the perturbers by their wakes and obtained fitting formulas that closely match the numerical solutions. The adiabatic sound speed is given by
@@
where $c_{\mathrm{s,0}}$ is the sound speed in the unperturbed medium. We shall see below that the subsonic regime, with $\mathcal{M}_{\mathrm{p}}<1$, is most relevant for this work. In this regime the Kim+08 fitting formulas are
@@
An analytical solution has been derived by Desjacques+22 that agrees with the numerical solutions of Kim+08, but as the analytical solution is cumbersome we choose to compare our solutions with the fitting formulas (17) and (18). [^3] See also Stahler+10 for an alternative analytic solution to the problem.
The $z$-component of the torque on a single perturber in the rest frame of the CM of the perturbers is given by $(a/2)\mathcal{F}\mathcal{I}_{\phi}$, where $\mathcal{F}$ is given by equation (12) and $\mathcal{I}_{\phi}$ by equation (14). Thus, for equal mass perturbers with equal Mach numbers, the torque on the binary is given by
@@
Equations (10) and (22) represent the approximate models for the torque on the binary that we will use to compare with the simulation.